OpenStudy (anonymous):

how can i solve n-6m=8 for n

OpenStudy (anonymous):

If you want to solve for n, how would you eliminate the -6m?

OpenStudy (anonymous):

by doing the opposite of subtracting?

OpenStudy (anonymous):

agree. so you would do what to both sides?

OpenStudy (anonymous):

adding

OpenStudy (anonymous):

so what do you get when you add 6m to BOTH sides of the equation?

OpenStudy (anonymous):

16?

OpenStudy (anonymous):

14?

OpenStudy (anonymous):

Dont you get 8 + 6m on the right side?

OpenStudy (anonymous):

yes so its 14m?

OpenStudy (anonymous):

You cannot combine 8 and 6m and say it is 14m. Because they are not like terms. You can add m's and m's...4m + 5m = 9m; you can add numbers together, 4+6 = 10, but you cannot add 6+ 8m and get 14m. So the final answer is n = 8 + 6m or n = 6m + 8

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

welcome.